# On the Duality Between Sharpness-Aware Minimization and Adversarial Training

### v1

@@ Abstract
Adversarial Training (AT), which adversarially perturb the input samples during training, has been acknowledged as one of the most effective defenses against adversarial attacks, yet suffers from a fundamental tradeoff that inevitably decreases clean accuracy. Instead of perturbing the samples, Sharpness-Aware Minimization (SAM) perturbs the model weights during training to find a more flat loss landscape and improve generalization. However, as SAM is designed for better clean accuracy, its effectiveness in enhancing adversarial robustness remains unexplored. In this work, considering the duality between SAM and AT, we investigate the adversarial robustness derived from SAM. Intriguingly, we find that using SAM alone can improve adversarial robustness. To understand this unexpected property of SAM, we first provide empirical and theoretical insights into how SAM can implicitly learn more robust features, and conduct comprehensive experiments to show that SAM can improve adversarial robustness notably without sacrificing any clean accuracy, shedding light on the potential of SAM to be a substitute for AT when accuracy comes at a higher priority. Code is available at https://github.com/weizeming/SAM_AT.
@@
Yihao Zhang1*1*1 *Hangzhou He1*1*1 *Jingyu Zhu1,2*1,2*1 , 2 *Huanran Chen3333Yifei Wang4444Zeming Wei11$\dagger$1 †
$1$Peking University $2$UC Berkeley $3$Beijing Institute of Technology $4$MIT
footnotetext: $*{}^{*}$qual contribution.
${}^{\dagger}$orrespondence to: Zeming Wei (weizeming@stu.pku.edu.cn).
Preprint.
@@ 1 Introduction
The existence of adversarial examples (Goodfellow et al., 2014; Szegedy et al., 2013) has raised serious safety concerns on the deployment of deep neural networks (DNNs). To defend against attacks of crafting adversarial examples, a variety of defense methods (Papernot et al., 2016; Xie et al., 2019) has been proposed. Adversarial training (AT) (Madry et al., 2017), which adds adversarial perturbations to the training samples in the training loop, has been acknowledged as one of the most effective adversarial defense paradigms (Athalye et al., 2018). Despite this success, AT suffers from an intrinsic limitation that decreases the clean accuracy, leading to a fundamental trade-off between accuracy and robustness (Tsipras et al., 2018; Zhang et al., 2019). A critical drawback of AT is that the perturbation of training samples makes the training distribution deviate from the natural data distribution.
Instead, Sharpness-Aware Minimization (SAM) (Foret et al., 2020) perturbs model weights yet keeps using the original samples during training. SAM is a novel training framework that regularizes the sharpness of the loss landscape by weight perturbation to improve generalization ability. So far, SAM has achieved remarkable success in modern machine learning (Bahri et al., 2021; Andriushchenko and Flammarion, 2022; Chen et al., 2023b). Many follow-up variants have been proposed (Kwon et al., 2021; Du et al., 2021; Chen et al., 2022). However, as SAM was originally designed for better natural generalization, its impact on adversarial robustness remains unexplored.
In this paper, considering the duality between SAM and AT, i.e. perturbing samples (AT) and perturbing weights (SAM), we explore whether SAM can also achieve better adversarial robustness. To be specific, if a neural network is robust against weight perturbation, it may also be robust against input perturbation due to the arithmetical combination of input and parameters to get the output. Surprisingly, we find that using SAM alone can notably improve adversarial robustness compared to standard optimization methods like Adam, which is an unexpected benefit of SAM. As illustrated in Table 1, SAM can achieve both higher robust and clean accuracy than standard training using Adam optimizer. Compared to the improvement in clean accuracy, the enhancement of robust accuracy (evaluated by AutoAttack (Croce and Hein, 2020b)) is significantly prominent (from 3.4% to 25.4%). By contrast, though AT can exhibit better robustness, their decrease in natural accuracy is unbearable in practical deployment. Motivated by this intriguing result, in this paper, we raise the two following questions and attempt to answer them: (1) Why does SAM improve adversarial robustness, and (2) To what extent SAM can be used as a substitution for AT.
@@
To study the two questions above, we first provide an intuitive understanding of the effect that SAM can have on improving adversarial robustness without sacrificing natural accuracy. Specifically, we show that both the robustness against the input and weight perturbation can enhance the stability of output, and such robustness on weight perturbation can help the model implicitly select the robust features (Tsipras et al., 2018). To verify this notion, we also provide theoretical evidence to support our intuitive understanding. Following the popular data distribution based on robust and non-robust features decomposition (Tsipras et al., 2018), we show that both SAM and AT can improve the robustness of the trained models by biasing more weight on robust features. To sum up, we answer the raised question (1) by concluding that SAM can improve adversarial robustness by implicitly learning robust features.
Furthermore, we conduct extensive experiments across different tasks (classification and segmentation) and data modalities (vision and text) to evaluate the robustness improvement of using SAM. Intriguingly, under various settings, SAM exhibits consistently better robustness than standard training methods while still maintaining higher natural performance. Despite AT can achieve higher robustness than SAM, it inevitably sacrifices natural performance which may be unaffordable for real-world applications. Besides, SAM also outperforms vanilla AT in terms of computational efficiency, since it only requires 1 additional step forward-backward pass than standard training. Though there are variants of AT designed for reducing computational cost (Shafahi et al., 2019), there are also efficient SAM variants proposed (Liu et al., 2022). Therefore, we answer the raised question (2) with the conclusion that SAM can be considered a lightweight substitute for AT under the following requirements: (i) no loss of any natural accuracy, and (ii) no significant increase in computational cost.
@@
    We provide both empirical and theoretical explanations to interpret how SAM can enhance adversarial robustness by showing that both input and weight perturbation can encourage the model to learn robust features.
@@
    We conduct comprehensive experiments on different tasks, data modalities, and model architectures on benchmark datasets to show the effectiveness of SAM in terms of enhancing robustness without sacrificing natural performance. We also suggest that SAM can be considered a lightweight substitute for AT under certain requirements.
@@ 2 Background and related work
2.1 Sharpness awareness minimization (SAM)
@@
where $\mathcal{D}$ is the data distribution, $L$ is the loss function, $w$ is the parameters of the model, $\|w\|_{2}^{2}$ is the regularization term and $\rho$ controls the magnitude of weight perturbation. To solve the inner maximization process, one-step gradient descent is applied. So far, SAM has become a powerful tool for enhancing the natural accuracy performance of machine learning models.
There are also many applications of SAM in other fields of research like language models (Bahri et al., 2021) and decentralized SGD (Zhu et al., 2023), showing the scalability of SAM to various domains. In addition, many improvements of the algorithm SAM spring up, like Adaptive SAM (ASAM) (Kwon et al., 2021), Efficient SAM (ESAM) (Du et al., 2021), LookSAM (Liu et al., 2022), Sparse SAM (SSAM) (Mi et al., 2022), Fisher SAM (Kim et al., 2022) and SAM-ON (Mueller et al., 2023), which add some modifications on SAM and further improve the generalization ability of the model. However, while these existing works focus on the natural generalization goal, the effectiveness of SAM on adversarial robustness remains unexplored.
2.2 Adversarial robustness
The adversarial robustness and adversarial training have become popular research topics since the discovery of adversarial examples (Szegedy et al., 2013; Goodfellow et al., 2014), which uncovers that DNNs can be easily fooled to make wrong decisions by adversarial examples that are crafted by adding small perturbations to normal examples. The malicious adversaries can conduct adversarial attacks by crafting adversarial examples, which cause serious safety concerns regarding the deployment of DNNs. So far, numerous defense approaches have been proposed (Xie et al., 2019; Bai et al., 2019; Mo et al., 2022; Chen et al., 2023a, 2024), among which adversarial training (AT) (Madry et al., 2017) has been considered the most promising defense method against adversarial attacks. AT can be formulated as the following optimization problem:
@@ 2.2 Adversarial robustness
where $\mathcal{D}$ is the data distribution, $\epsilon$ is the margin of perturbation, $w$ is the parameters of the model and $L$ is the loss function (e.g. the cross-entropy loss). For the inner maximization process, the Projected Gradient Descent (PGD) attack is commonly used to generate the adversarial example:
@@
Though improves adversarial robustness effectively, adversarial training has exposed several defects such as computational overhead (Shafahi et al., 2019), class-wise fairness (Xu et al., 2021; Wei et al., 2023b) and robust overfitting (Rice et al., 2020; Wang et al., 2023; Wei et al., 2023a) among which the decreased natural accuracy (Tsipras et al., 2018; Zhang et al., 2019) has become the major concern.
In the context of adversarial robustness, several works also attempt to introduce a flat loss landscape in adversarial training (Wu et al., 2020; Yu et al., 2022a, b). The most representative one is Adversarial Weight Perturbation (AWP) (Wu et al., 2020), which simultaneously adds perturbation on examples and feature space to apply SAM and AT. However, AWP also suffers from a decrease in natural accuracy which is even lower than AT in some cases. We assume that this is because both perturbing the inputs and parameters significantly raise the difficulty of robust learning. Besides this thread of work, the adversarial robustness derived from SAM alone has not been explored.
@@ 3 Empirical understanding
In this section, we provide an intuitive understanding of the duality between SAM and AT, which explains how SAM improves adversarial robustness. We start by rewriting the optimization objective of SAM and AT in a unified form and omit the regularization term $\lambda\|w\|_{2}^{2}$:
@@
To be more specific, we illustrate our understanding with a middle linear layer in a model, which extracts feature $z$ from input $x$: $z=Wx$. During AT, we add perturbations directly to the input space, resulting in $x\leftarrow x+\delta$. However, in SAM, the perturbation is not directly applied to the input space, but to the parameter space as $W\leftarrow W+\delta$. This leads to $Wx+W\delta$ for input perturbation and $Wx+\delta x$ for parameter perturbation. If the weight $W$ can keep $Wx$ more robust against small perturbations around it in the subsequent layers, it will be also beneficial to improve sample robustness around $x$.
In addition, we discuss the attack (perturbation) strength of AT and SAM. For SAM, the perturbation is relatively more moderate, as its perturbations are conducted in the weight space and do not change the original input. On the other hand, in order to achieve the best robustness by eliminating the non-robust features (Ilyas et al., 2019), AT applies larger and more straightforward perturbations to the input space, leading to better robustness but a loss in natural accuracy.In summary, SAM applies small perturbations implicitly to the feature space to maintain good natural accuracy performance, while AT utilizes direct input perturbations, which may result in a severe loss in natural accuracy. We provide more theoretical evidence to support these claims in the next section.
4 Theoretical analysis
@@ 4.1 A binary classification model
to measure the weight of robust features involved in the model prediction.
@@ 4.3 Adversarial training (AT)
Now let’s consider when AT is applied. Recall that in this case, the loss function is no longer the standard one but the expected adversarial loss
@@
Similar to standard training, we can derive the robust feature weight from: which can be stated in the following theorem:
@@ 4.4 Sharpness-Aware Minimization (SAM)
From Theorem 4.2 and 4.3 we can see that both $W_{R}^{AT}$ and $W_{R}^{SAM}$ are greater than $W_{R}^{*}$, which indicates both SAM and AT encourages the trained model to learn more robust features. However, the qualitative relation is not sufficient to quantify how much robustness SAM achieves compared to adversarial training, and we attempt to step further by quantitatively estimating the $W_{R}^{SAM}$ in the following theorem:
@@
In the classification problem above, suppose that $\epsilon$ is small, we have
@@ 4.5 Relation between SAM and AT
All proofs can be found in Appendix A. To sum up, we can conclude that AT utilizes explicit and direct perturbations for eliminating non-robust features, while SAM leverages implicit and moderate perturbations to learn robust features. This is consistent with our empirical understanding in Section 3 and we also verify these claims with experiments in the following section.

### v2

@@ Abstract
Adversarial Training (AT), which adversarially perturb the input samples during training, has been acknowledged as one of the most effective defenses against adversarial attacks, yet suffers from inevitably decreased clean accuracy. Instead of perturbing the samples, Sharpness-Aware Minimization (SAM) perturbs the model weights during training to find a more flat loss landscape and improve generalization. However, as SAM is designed for better clean accuracy, its effectiveness in enhancing adversarial robustness remains unexplored. In this work, considering the duality between SAM and AT, we investigate the adversarial robustness derived from SAM. Intriguingly, we find that using SAM alone can improve adversarial robustness. To understand this unexpected property of SAM, we first provide empirical and theoretical insights into how SAM can implicitly learn more robust features, and conduct comprehensive experiments to show that SAM can improve adversarial robustness notably without sacrificing any clean accuracy, shedding light on the potential of SAM to be a substitute for AT when accuracy comes at a higher priority. Code is available at https://github.com/weizeming/SAM_AT.
@@ 1 Introduction
The existence of adversarial examples (Goodfellow et al., 2014; Szegedy et al., 2013) has raised serious safety concerns on the deployment of deep neural networks (DNNs). To defend against attacks of crafting adversarial examples, a variety of defense methods (Papernot et al., 2016; Xie et al., 2019; Cohen et al., 2019; Chen et al., 2023a) has been proposed. Adversarial training (AT) (Madry et al., 2017), which adds adversarial perturbations to the training samples in the training loop, has been acknowledged as one of the most effective adversarial defense paradigms (Athalye et al., 2018). Despite this success, AT suffers from an intrinsic limitation that decreases the clean accuracy, leading to a fundamental trade-off between accuracy and robustness (Tsipras et al., 2018; Zhang et al., 2019). A critical drawback of AT is that the perturbation of training samples makes the training distribution deviate from the natural data distribution.
Instead, Sharpness-Aware Minimization (SAM) (Foret et al., 2020) perturbs model weights yet keeps using the original samples during training. SAM is a novel training framework that regularizes the sharpness of the loss landscape by weight perturbation to improve generalization ability. So far, SAM has achieved remarkable success in modern machine learning (Bahri et al., 2021; Andriushchenko and Flammarion, 2022; Chen et al., 2023b) and many of its follow-up variants have been proposed (Kwon et al., 2021; Du et al., 2021; Chen et al., 2022). However, as SAM was originally designed for better natural generalization, its impact on adversarial robustness remains unexplored.
In this paper, considering the duality between SAM and AT, i.e. perturbing samples (AT) and perturbing weights (SAM), we explore whether SAM can also achieve better adversarial robustness. Surprisingly, we find that using SAM alone can notably improve adversarial robustness compared to standard optimization methods like Adam, which is an unexpected benefit of SAM. As illustrated in Table 1, SAM can achieve both higher robust and clean accuracy than standard training. Compared to the improvement in clean accuracy, the enhancement of robust accuracy (evaluated by AutoAttack (Croce and Hein, 2020b)) is significantly prominent (from 3.4% to 25.4%). By contrast, though AT can exhibit better robustness, their decrease in natural accuracy is unaffordable in practical deployment. Motivated by this intriguing observation, we raise the two following questions and attempt to answer them:
(1) Why can SAM improve adversarial robustness, and
(2) Under what condition can SAM be deployed as a substitution for AT.
@@
To study the two questions above, we first provide an intuitive understanding of why SAM can improve adversarial robustness without sacrificing natural accuracy. Specifically, we show that weight perturbation during training can help the model implicitly learn the robust features (Tsipras et al., 2018). To verify this notion, we also provide theoretical insights to support our intuitive understanding. Following the popular data distribution based on robust and non-robust features decomposition (Tsipras et al., 2018), we show that both SAM and AT can improve the robustness of the trained models by biasing more weight on robust features. To sum up, we answer the raised question (1) by concluding that SAM can improve adversarial robustness by implicitly learning robust features.
Furthermore, we conduct extensive experiments across different tasks, data modalities, and model architectures to evaluate the robustness improvement of using SAM, where we show that under various settings, SAM exhibits consistently better robustness than standard training methods while still maintaining higher natural performance. Besides, though AT can achieve higher robustness than SAM, it inevitably sacrifices natural performance which may be unaffordable for real-world applications. Therefore, we answer the raised question (2) with the conclusion that SAM can be used as a substitute for AT when accuracy is more important but better robustness is preferred.
@@
    We provide empirical and theoretical insights to understand how SAM can enhance adversarial robustness by showing that both input and weight perturbation can encourage the model to learn robust features.
@@
    We conduct extensive experiments to show the effectiveness of SAM in terms of enhancing robustness without sacrificing natural performance. We also suggest that SAM can be considered a lightweight substitute for AT when accuracy comes at a higher priority.
@@ 2 Background and Related Work
2.1 Sharpnes-Aware Minimization (SAM)
@@
where $\mathcal{D}$ is the data distribution, $L$ is the loss function, $w$ is the parameters of the model, $\|w\|_{2}^{2}$ is the regularization term and $\rho$ controls the magnitude of weight perturbation. To solve the inner maximization process, one-step gradient descent is applied.
So far, SAM has become a powerful tool for enhancing the natural accuracy performance of machine learning models. There are also many applications of SAM in other fields of research like language models (Bahri et al., 2021) and decentralized SGD (Zhu et al., 2023), showing the scalability of SAM to various domains. In addition, many improvements of the algorithm SAM spring up, like Adaptive SAM (ASAM) (Kwon et al., 2021), Efficient SAM (ESAM) (Du et al., 2021), LookSAM (Liu et al., 2022), Sparse SAM (SSAM) (Mi et al., 2022), Fisher SAM (Kim et al., 2022) and SAM-ON (Mueller et al., 2023), which add some modifications on SAM and further improve the generalization ability of the model. However, while these existing works focus on the natural generalization goal, the effectiveness of SAM on adversarial robustness remains unexplored.
2.2 Adversarial Robustness and Adversarial Training
The adversarial robustness and adversarial training have become popular research topics since the discovery of adversarial examples (Szegedy et al., 2013; Goodfellow et al., 2014), which uncovers that DNNs can be easily fooled to make wrong decisions by adversarial examples that are crafted by adding small perturbations to normal examples. The malicious adversaries can conduct adversarial attacks by crafting adversarial examples, which cause serious safety concerns regarding the deployment of DNNs. So far, numerous defense approaches have been proposed (Xie et al., 2019; Bai et al., 2019; Cohen et al., 2019; Chen et al., 2024), among which Adversarial Training (AT(Madry et al., 2017) has been considered the most promising defense method against adversarial attacks. AT can be formulated as the following optimization problem:
@@ 2.2 Adversarial Robustness and Adversarial Training
where $\mathcal{D}$ is the data distribution, $\epsilon$ is the margin of perturbation, $w$ is the parameters of the model and $L$ is the loss function (e.g. the cross-entropy loss). For the inner maximization process, the Projected Gradient Descent (PGD) (Madry et al., 2017) attack is commonly used to generate the adversarial example:
@@
Though improves adversarial robustness effectively, adversarial training has exposed several defects such as computational overhead (Shafahi et al., 2019), class-wise fairness (Xu et al., 2021; Wei et al., 2023b) and robust overfitting (Rice et al., 2020; Wang et al., 2023; Wei et al., 2023a) among which the decreased natural accuracy (Tsipras et al., 2018; Zhang et al., 2019) has become the major concern. One explanation for this drawback of AT is perturbing the samples during training leads the training sample distribution to deviate from the natural data (Ilyas et al., 2019).
In the context of adversarial robustness, several works also attempt to introduce a flat loss landscape in adversarial training (Wu et al., 2020; Yu et al., 2022a, b). The most representative one is Adversarial Weight Perturbation (AWP) (Wu et al., 2020), which simultaneously adds perturbation on examples and feature space to apply SAM and AT. However, AWP also suffers from a decrease in natural accuracy which is even lower than AT in some cases, which we assume is because perturbing both the inputs and parameters significantly raises the difficulty of robust learning. We compare SAM and AWP in our experiment section. Besides this thread of work, the adversarial robustness derived from SAM alone has not been explored.
@@ 3 Empirical Understanding
In this section, we provide an intuitive explanation to empirically understand how SAM improves adversarial robustness by demonstrating the duality between SAM and AT. Specifically, considering the arithmetical duality of the input and parameters to get the output in a specific layer, we can assume that the robustness against weight perturbation may also lead to robustness against input perturbation.
We start by rewriting the optimization objective of SAM and AT in a unified form and omit the regularization term $\lambda\|w\|_{2}^{2}$ as follows:
@@
To be more formalized, we illustrate our understanding with a middle linear layer in a model, which extracts feature $z$ from input $x$: $z=Wx$. During AT, we add perturbations directly to the input space, resulting in $x\leftarrow x+\delta$. However, in SAM, the perturbation is not directly applied to the input space, but to the parameter space as $W\leftarrow W+\delta$. This leads to $Wx+W\delta$ for input perturbation and $Wx+\delta x$ for parameter perturbation. If the weight $W$ can keep $Wx$ more robust against small perturbations around it in the subsequent layers, it will be also beneficial to improve sample robustness around $x$.
Besides, we discuss the attack (perturbation) strength of AT and SAM. For SAM, the perturbation is relatively more moderate, as its perturbations are conducted in the weight space and do not change the original input. On the other hand, in order to achieve the best robustness by eliminating the non-robust features (Ilyas et al., 2019), AT applies larger and more straightforward perturbations to the input space, leading to better robustness but a loss in natural accuracy, which is not the original goal of SAM, but it is the goal of AT.
In summary, our intuitive analysis suggests that SAM applies small perturbations implicitly to the feature space to maintain good natural accuracy performance, while AT utilizes direct input perturbations, which may result in a severe loss in natural accuracy. We provide more theoretical evidence to support these claims in the next section.
4 Theoretical Insights
@@ 4.1 A Binary Classification Model
to measure the weight of robust features involved in the model prediction. We use the following lemma to justify the fundamental relationship between $W_{R}$ and the adversarial robustness of the model:
Theorem 4.1.
The robust accuracy ($R_{A}$) of this model, defined as
$R_{A}=\mathbb{E}_{\mathbf{x},y\sim D}\underset{||\delta||<\epsilon}{\min}\mathbf{1}\{{f_{\boldsymbol{w}}(x+\delta)=y}\}.$ (8)
is a monotonic increasing function of $W_{R}$ under condition $\epsilon<\eta$ and $0<W_{R}<W_{R}^{AT}$(defined in (11)).
@@ 4.3 Adversarial Training (AT)
Now we consider when AT is applied to train the model. Recall that in this case, the loss function is no longer the standard one but the expected adversarial loss
@@
Similar to standard training, we can derive the robust feature weight from the following theorem:
@@ 4.4 Sharpness-Aware Minimization (SAM)
From Theorem 4.3 and 4.4 we can see that both $W_{R}^{AT}$ and $W_{R}^{SAM}$ are greater than $W_{R}^{*}$, which indicates both SAM and AT encourages the trained model to learn more robust features. However, the qualitative relation is not sufficient to quantify how much robustness SAM achieves compared to adversarial training, and we attempt to step further by quantitatively estimating the $W_{R}^{SAM}$ in the following theorem:
@@
In the classification problem above, suppose that $\epsilon>0$ is small, we have
@@ 4.5 Relation between SAM and AT
All proofs can be found in Appendix A. While we acknowledge that $W_{R}$ is only analyzed under a simple model and cannot be directly generalized to multiple-layer networks, we believe the insights delivered from $W_{R}$ can be generalized to DNNs. To sum up, we can conclude that AT utilizes explicit and direct perturbations for eliminating non-robust features, while SAM leverages implicit and moderate perturbations to learn robust features. This is consistent with our empirical understanding in Section 3 and we also verify these claims with experiments in the following section.
@@ 5 Experiment
5.1 Image Classification
5.1.1 Experimental settings
Datasets and models. We examine the robustness of SAM on CIFAR-{10,100} (Krizhevsky et al., 2009) and TinyImageNet (Chrabaszcz et al., 2017) datasets. We mainly conduct our experiment with PreActResNet-18 (PRN-18) (He et al., 2016b). To demonstrate the scalability of SAM, we also include the Wider ResNet (WRN-28-10) (Zagoruyko and Komodakis, 2016) and the vision transformer architecture (DeiT)(Touvron et al., 2021). To evaluate the corruption robustness of various models, we also use the CIFAR-10C dataset (Hendrycks and Dietterich, 2019), a common corruptions dataset for CIFAR10 with 1000 randomly selected examples from CIFAR-10C and level-3 of perturbation severity for each corruption types (fog, snow, etc).
Attacks for robustness evaluations. We assess the model’s resistance through a multifaceted approach to thoroughly explore its robustness against different threat models. We apply the Fast Gradient Sign Method (FGSM(Goodfellow et al., 2015) with a perturbation bound of $\ell_{\infty}$-$\epsilon=1/255$, and a 10-step Projected Gradient Descent (PGD) attack under $\ell_{\infty}$ and $\ell_{2}$ norms with bounds of $\epsilon=1/255$ and $32/255$, respectively. Moreover, we incorporate the AutoAttack (AA.) (Croce and Hein, 2021) to ensure a reliable evaluation. Besides, we also consider other popular attacks like StAdv (Xiao et al., 2018), FAB (Croce and Hein, 2020a) (10 steps under $\ell_{2}$-$\epsilon=32/255$). For black-box assessment, we utilize Pixle (Pomponi et al., 2022), a strong pixel-rearranging black-box attack, constrained to a maximum of five iterations. The attacks are mostly implemented by the torchattacks (Kim, 2020) framework to ensure a reliable assessment. For corruption robustness (Hendrycks and Dietterich, 2019), which measures the robustness of DNNs against a wide range of real-world disturbances like noise, blur, or weather variations, we evaluate models on RobustBench (Croce et al., 2021), a standardized benchmark designed to evaluate the general robustness. We ran all experiments three times independently to report the average result and omitted the standard deviations since they are small (less than 0.5%) and do not affect our claims.
5.1.2 Experimental results
Comparison with standard training. As shown in Tables 2, 3 and 4, all models trained with SAM exhibit significantly better natural accuracy and robustness compared to those trained with standard training (ST). In particular, higher robustness is achieved by using larger values of $\rho$ with SAM. Taking the CIFAR-100 dataset as an example, the model trained with $\rho=0.4$ demonstrates even multiple robust accuracy than ST, and its natural accuracy is still higher than that of ST. Compared to the improvement in natural accuracy (approximately 2%), the increase of robustness is more significant (more than 10% in average). Moreover, on large datasets like Tiny-ImageNet, SAM still surpasses ST by a large margin. Therefore, we conclude that SAM with a relatively larger weight perturbation bound $\rho$ is a promising technique for enhancing adversarial robustness without sacrificing natural accuracy. Notably, our results are consistent with concurrent work showing that the robustness of SGD is better than Adam (Ma et al., 2023).
Table 4: Comparison on Tiny ImageNet.
Config Natural AA. StAdv
SGD 57.4 2.5 3.4
SAM ($\rho=0.4$) 57.9 10.3 13.3
SAM ($\rho=0.5$) 57.7 10.4 12.7
$\ell_{\infty}$-AT($\epsilon=8/255$) 32.3 27.7 25.1
$\ell_{2}$-AT($\epsilon=128/255$) 41.6 32.8 30.2